OpenStudy (anonymous):

the equation of a circle when endpoints of a diameter at (-2,-9) and (0,-5)

OpenStudy (anonymous):

when the ends of the diameter are given as(x1,y1) and (x2,y2) the eq of circle is given by (x-x1)(x-x2) +(y-y1)(y-y2)=0 i hope rest u will be able to do

OpenStudy (anonymous):

idk what the heck u just did

OpenStudy (anonymous):

the eq of the circle is given by (x+2)(x-0)+(y+9)(y+5)=0 or x^2+2x+y^2+14y+45=0 or x^+y^2+2x+14y+45=0